Circular No. 7802

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE Further to IAUC 7801, B. Barris reports the following additional ap-

parent supernovae. The tabulated redshift for SN 2002ab is that of the type-Ia supernova itself. SN 2002Y and SN 2002Z are not present in images taken on 2001 Oct. 22 (limiting mag 26.0); the remaining objects are not present in images taken on 2001 Nov. 18 and 19 (limiting mag 25.8).

SN  2002 UT  α_{2000}  δ_{2000}  I_{1}  I_{2}  z
2002Y Jan. 9 8^{h}48^{m}50.^{s}21 +44^{o}13′03′′.0 24.2 24.5 0.56
2002Z Jan. 9 8 49 32.97 +44 03 27.6 24.2 24.3 0.395
2002aa Jan. 9 7 48 45.28 +10 18 00.8 24.3 24.2 0.95
2002ab Jan. 9 7 48 55.70 +10 06 06.3 22.5 22.5 0.42
2002ac Jan. 9 7 49 39.61 +10 19 08.2 23.9 23.8
2002ad Jan. 9 10 50 12.19 +57 31 11.6 23.0 23.2 0.52
2002ae Jan. 9 10 50 51.02 +57 09 04.6 23.8 23.6
2002af Jan. 9 10 52 29.89 +57 39 44.0 24.6 24.1
2002ag Jan. 9 10 53 12.71 +57 00 25.9 23.5 24.2
2002ah Jan. 9 10 53 19.50 +57 00 36.6 24.2 24.4
2002ai Jan. 9 10 54 21.27 +57 27 58.2 23.4 23.3

SUPERNOVA 2002D IN NGC 594 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002D (cf. IAUC 7797), obtained by M. Calkins on Jan. 21.11 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. P-Cyg lines of hydrogen are apparent on a blue continuum. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 5413 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 8500 km/s.

SUPERNOVA 2002G IN CGCG 189-024 Matheson et al. also report that a spectrum of SN 2002G (cf. IAUC

7797), obtained by Calkins as above on Jan. 20.52 UT, shows it to be a type-Ia supernova before maximum. Adopting the NED recession velocity of 10 114 km/s for the host galaxy, the supernova expansion velocity is ∼ 12 300 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (cf. IAUC 7774) is 7 ± 2 days before maximum light.

2002 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green